OpenStudy (anonymous):

what is the simplified form of 11sqrt(8)+6sqrt(12)-5sqrt(2)

OpenStudy (campbell_st):

you need to write each radical as a multiple of sqrt(2) \[\sqrt{8} = \sqrt{4\times2} = \sqrt{4}\times \sqrt{2} = 2\sqrt{2}\] so \[11\sqrt{8} = 11\times2\sqrt{2} = 22\sqrt{2}\] you need to do the same type of thing for \[6\sqrt{12}\]

OpenStudy (campbell_st):

not quite.... here is the start \[6\sqrt{12} = 6\sqrt{4 \times3}\] so in your question you will have 2 like terms, (sqrt(2) and one that is sqrt(3)

OpenStudy (campbell_st):

so at the moment you have \[22\sqrt{2} + 6\sqrt{4\times3} - 5\sqrt{2}\]

OpenStudy (anonymous):

okay

OpenStudy (campbell_st):

then you have \[(22 - 5)\sqrt{2} + 6\sqrt{4\times3}\]

OpenStudy (anonymous):

okay

OpenStudy (campbell_st):

\[17\sqrt{2} + 6\sqrt{4}\times \sqrt{3}\] I'll let you finish it

OpenStudy (anonymous):

well what would i do because there's no like terms?

OpenStudy (campbell_st):

\[17\sqrt{2} + 12\sqrt{3}\]